OpenStudy (anonymous):

Use the quadratic formula to solve x^2 + 9x + 10 = 0. Estimate irrational solutions to the nearest tenth.

OpenStudy (anonymous):

@iambatman

OpenStudy (kainui):

Where are you having trouble using the quadratic formula @DragonFlame ?

OpenStudy (anonymous):

-6.6 -1.4

OpenStudy (anonymous):

you just need to use \[-b \pm \sqrt{b^2-4ac}\div2a\]

OpenStudy (anonymous):

\[\huge x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

a=1 b=9 c=10

OpenStudy (anonymous):

@Kainui I don't know how to put abc into the formula

OpenStudy (anonymous):

so you get\[-9\pm \sqrt{81-4(10)}\div \]

OpenStudy (anonymous):

2a

OpenStudy (anonymous):

So just / by 2?

OpenStudy (anonymous):

-9+-sqrt41/2?

OpenStudy (anonymous):

so then u solve it